OpenStudy (anonymous):

Suppose an astronaut drops a feather from 1.2 m above the surface of the Moon. If the acceleration due to gravity on the Moon is 1.62 m/s2 downward, how long does it take the feather to hit the Moon's surface? s

OpenStudy (anonymous):

use the kinematic equation: s = s + vt + (1/2)at^2 plug in what we know: 0 = 1.2 + 0 + (1/2)(-1.62)t^2 solve for t

OpenStudy (anonymous):

ohh i gotacha thank you!!

OpenStudy (anonymous):

hes wrong

OpenStudy (anonymous):

@zachy1234 what would you suggest as a solution?